The coherence length for sodium D₂ line is 2.5cm. To deduce the coherence time and spectral width of the line. Given wavelength of D₂ line=6000 angstrom.
1
Expert's answer
2019-11-26T10:36:04-0500

We can writer

"l=tc (1)"

where l is the coherence length, t is the coherence time, c is velocity of light in vacuum

Using (1) we get

t=8.3×10-10 s

The spectral width of the line is given by formula

"Δλ=\frac {λ×λ}{l} (2)"

Using (2) we get

Δλ=14.4×10-12 m


Answer

8.3×10-10 s

14.4×10-12 m